OpenStudy (anonymous):

In india, water lilies grow extremely fast. In one pond, a lily grew so fast that each day it doubled in the area it covered. In 26 days it covered the pond. How long would it take 2 lilies to cover the pond?

OpenStudy (anonymous):

it takes one doubling less when we start at lillies=2 instead of lillies=1 right? we omit the day where 1 turns to 2 25 days

OpenStudy (anonymous):

how would i write it out?

OpenStudy (anonymous):

given: maxlillies = 1 * (2)^26 asked: maxlillies = 2 * (2)^x